OpenStudy (anonymous):

George has found a relationship between the height of a person and his bowling score. The table below shows the data collected by George: Height cm 152 160 164 170 180 162 154 158 168 Bowling Score 51 55 57 60 65 55 52 54 59 Part A: What would most likely be the bowling score of a person who has a height of 156 cm? Part B: Predict a possible correlation coefficient for the data in the table and explain why you think your prediction is a good value for the data. Part C: George says that if the bowling ball is rolled with a greater force, the bowling score will be higher. Is this an example of correlation or causation?

OpenStudy (anonymous):

@myininaya @trexnapora @kawaiicat123 @KlOwNlOvE

OpenStudy (anonymous):

@Gebooors

OpenStudy (anonymous):

look at score for 152 and 160 whats the difference

OpenStudy (anonymous):

160 is higher the 152

OpenStudy (anonymous):

by how much?

OpenStudy (anonymous):

55-51=

OpenStudy (anonymous):

\[\frac{ y2-y1 }{ x2-x1 }\]

OpenStudy (anonymous):

\[\frac{ 55-51 }{ 160-152 }\]

OpenStudy (anonymous):

oh wait im so sorry its 53 isn't it?

OpenStudy (anonymous):

@KlOwNlOvE ??

OpenStudy (anonymous):

sorry forgot i forgot i was here can you do that equation for me